OpenStudy (anonymous):

The intake pipe can fill a certain tank in 6 hours when the outlet pipe is closed, but with the outlet pipe open, it takes 9 hours. How long would it take the outlet pipe to empty a full tank?

OpenStudy (anonymous):

Full 1 = In X hours or F = I(6) and Full 2 = In X hours - Out X hours or F = I(9) - O(9) so setting them equal to each other since they both mean the tank is full I(6) = I(9) - O(9) rearrange I(6) - I(9) = - O(9) combining like values - I(3) = -O(9) multiply both sides by negative one I(3) = O(9) since I(6) equals a full tank we need to double the equation I(6) = O(18) So a full tank will drain out in 18 hours